# Nested ensemble Kalman filter for static parameter inference in nonlinear state-space models

### v1

@@ 3.2 Ensemble MCMC
where $\psi_{T}^{(\theta)}(\tilde{x}_{0:T}^{1:N},x_{0:T}^{1:N})$ denotes the joint density of the entire ensemble $\{\tilde{x}_{0:T}^{1:N},x_{0:T}^{1:N}\}$. Although (14) is typically intractable, it can be sampled using MCMC, and the parameter samples retained to give (dependent) draws from $\widehat{\pi}_{\textrm{enkf},T}(\theta|y_{0:T})$, the $\theta$-marginal of (14). The resulting approach, termed ensemble MCMC (eMCMC) is detailed in drovandi22 (see also Stroud2018; Katzfuss2019). In brief, at iteration $k$, a new $\theta^{*}$ is proposed from a kernel $q(\theta^{k-1},\cdot)$. Then, sampling of the corresponding ensemble $\{\tilde{x}_{0:T}^{*,1:N},x_{0:T}^{*,1:N}\}$ from $\psi_{T}^{(\theta^{*})}(\cdot,\cdot)$ and evaluation of the observed data likelihood $\widehat{p}_{\textrm{enkf},T}^{(\theta^{*})}(y_{0:T}|\tilde{x}_{0:T}^{*,1:N})$, is achieved through recursive application of Algorithm 3. The parameter $\theta^{*}$ is retained with probability
@@
When applied sequentially, eMCMC is wasteful in terms of both storage and computational costs; upon receipt of a new observation (say $y_{T+1}$), previous parameter samples are discarded and eMCMC is run from scratch. Nevertheless, we use eMCMC for benchmarking the accuracy of various sequential approaches in the context of static parameter inference, for example, when running PMMH would be computationally prohibitive. It is also a key component of the resample-move step described in Section 4.
@@ 3.3 State augmentation
The EnKF can be directly modified to simultaneously estimate both hidden states and static parameters by augmenting the state vector $x_{t}$ to incorporate $\theta$ (see e.g. Evensen2007; mitchell21). We write $z_{t}=(x_{t}^{\prime},\theta_{t}^{\prime})^{\prime}$ with $x_{t}$ evolving according to $p_{t}^{(\theta)}(\cdot|x_{t-1})$ and $\theta_{t}=\theta_{t-1}=\theta$. The observation model becomes $f_{t}^{(\theta)}(y_{t}|z_{t})=\mathcal{N}(y_{t};H_{z}z_{t},R)$ with $H_{z}=(H,0_{d_{y}\times d_{\theta}})$, where $0_{d_{y}\times d_{\theta}}$ is a $d_{y}\times d_{\theta}$ matrix of zeros and $H$ is not a function of $\theta$.
@@
To alleviate under-dispersion in the parameter ensemble, for example when small ensemble sizes are considered (e.g. Ruckstuhl18), the parameter vector can be allowed to dynamically evolve, for example using the artificial evolution kernel of liu2001, which gives
@@
where $\bar{\theta}_{t-1}$ and $V_{t-1}$ are the sample mean and variance of $\theta_{t-1}^{1},\ldots,\theta_{t-1}^{N}$. Setting $a=\sqrt{1-h^{2}}$ corrects for over-dispersion; the practitioner can then choose the smoothing parameter $h$, for example using the discounting method of liu2001. Step $t$ of the augmented EnKF is summarised by Algorithm 4.
@@ 3.4 Particle ensemble Kalman filter (PEnKF)
The augmented EnKF makes the strong assumption that the parameters and observations follow a linear Gaussian structure. An alternative approach for dealing with the static parameters is to marginalise out the hidden states in a sequential fashion. Recall the factorisation of $\pi_{t}(\theta|y_{0:t})$ in (5) and the assumption of a weighted sample of size $M$, $\{\theta^{1:M},u_{t-1}^{1:M}\}$, from $\pi_{t-1}(\theta|y_{0:t-1})$. Upon receipt of $y_{t}$, and in the context of time-varying parameters, Katzfuss2019 (see also frei11, among others) propose to use the observed-data likelihood under the EnKF to reweight parameter particles. We adapt the resulting particle ensemble Kalman filter (PEnKF) to our problem as follows.
@@ 4 Nested ensemble Kalman filter (NEnKF)
From the perspective of static parameter filtering, the EnKF-based methods described in Section 3 attempt to mitigate particle degeneracy via an artificial evolution kernel over the static parameters. This can be sensitive to the choice of tuning parameters (vieira2016). The SMC2 algorithm described in Section 2 provides a more principled approach to alleviating particle degeneracy: if triggered at some time $t$, a resample-move step mutates parameter particles through a (pseudo-marginal) Metropolis-Hastings kernel that has $\pi_{t}(\theta|y_{0:t})$ as its invariant distribution. However, both the weighting and resample-move step in SMC2 require running a particle filter over states, which can require a computationally prohibitive number of state particles $N$ in high-dimensional settings. For a specific example with $d_{x}=d_{y}=:n$, Katzfuss2019 show that the variance of the logarithm of observed-data likelihood under the EnKF, $\textrm{Var}[\log\widehat{p}_{\textrm{enkf},T}^{(\theta)}(y_{0:T}|\tilde{x}_{0:T}^{1:N})]$, can be controlled by scaling $N$ linearly with $n$, whereas $N$ should be scaled exponentially with $n$ when controlling $\textrm{Var}[\log\widehat{p}_{T}^{(\theta)}(y_{0:T}|x_{0:T}^{1:N},a_{0:T-1}^{1:N})]$, computed using a particle filter. When comparing the efficiency of ensemble MCMC and particle MCMC across several examples, drovandi22 find that the former can tolerate a much smaller value of $N$ compared to the latter.
@@ 4.1 NEnKF algorithm
${\alpha}_{\textrm{enkf}}(\theta^{*}|\theta^{i})=1\wedge\frac{\pi(\theta^{*})}{\pi(\theta^{i})}\times\frac{\widehat{p}_{\textrm{enkf},t}^{(\theta^{*})}(y_{0:t}|\tilde{x}_{0:t}^{*,1:N})}{\widehat{p}_{\textrm{enkf},t}^{(\theta^{i})}(y_{0:t}|\tilde{x}_{0:t}^{i,1:N})}\times\frac{q(\theta^{*},\theta^{i})}{q(\theta^{i},\theta^{*})}.$ (18)
@@
   2. Resample. Normalise the weights to give $u_{t}^{i}=\tilde{u}_{t}^{i}/\sum_{k=1}^{M}\tilde{u}_{t}^{k}$. Sample an index $b_{i}$ from $\{1,\ldots,M\}$ with probabilities $u_{t}^{1:M}$. Put $\{\theta^{i},u_{t}^{i}\}:=\{\theta_{t}^{b_{i}},1/M\}$ and $x_{t}^{i,1:N}:=x_{t}^{b_{i},1:N}$.
   3. Move / mutate. Propose $\theta^{*}\sim q(\theta^{i},\cdot)$. Perform iterations $1,\ldots,t$ of Algorithm 3 to obtain $\widehat{p}_{\textrm{enkf},t}^{(\theta^{*})}(y_{0:t}|\tilde{x}_{0:t}^{*,1:N})$ and $x_{t}^{*,1:N}$. With probability (18), put $\theta^{i}=\theta^{*}$, $\widehat{p}_{\textrm{enkf},t}^{(\theta^{i})}(y_{0:t}|\tilde{x}_{0:t}^{i,1:N})=\widehat{p}_{\textrm{enkf},t}^{(\theta^{*})}(y_{0:t}|\tilde{x}_{0:t}^{*,1:N})$ and $x_{t}^{i,1:N}=x_{t}^{*,1:N}$.
@@
Following the approach used in SMC2, we also consider a mechanism for dynamically updating the number of ensemble members $N$. After completion of the mutation step, if updating $N$ is triggered then for each particle $\theta^{i}$, ensemble members $x_{0:t}^{i,1:N}$ are replaced by $\check{x}_{0:t}^{i,1:\check{N}}$ using the generalised importance sampling strategy of delMoral06, with an artificial backward kernel chosen to give an incremental weight of 1 (botha23). Hence, for each particle $\theta^{i}$, a new ensemble $\check{x}_{0:t}^{i,1:\check{N}}$ is generated through repeated application of Algorithm 3. Once the updating of $N$ is triggered, we adapt the method of duan15 to our context, by setting $\check{N}=\hat{\sigma}_{N}^{2}N$, where $\hat{\sigma}_{N}^{2}$ is an estimate of the variance of $\widehat{p}_{\text{enkf},t}^{(\bar{\theta})}(y_{0:t}|\tilde{x}_{0:t}^{1:N})$, based on, say, $r$ runs of the EnKF with $\theta$ fixed at an estimate of some central posterior value (e.g. the posterior mean). The value of $\hat{\sigma}_{N}^{2}$ can also be used to trigger updating of $N$, for example, if $\hat{\sigma}_{N}^{2}>1.5$. We refer the reader to botha23 for further discussion, noting their critique that, in the SMC2 setting, this approach can be sensitive to the initial value of $N$. However, and as discussed at the beginning of this section, $\widehat{p}_{\text{enkf},t}^{(\theta)}(y_{0:t}|\tilde{x}_{0:t}^{1:N})$ is likely to be less sensitive to $N$ than the corresponding particle filter estimator.
@@ 4.2 Resample-move: delayed-acceptance kernel
As with SMC2, the most computationally expensive step of the NEnKF algorithm is the resample-move step; if executed at time $t$, the cost is $\mathcal{O}(tMN)$. We, therefore, aim to avoid expensive calculation of the likelihood $\widehat{p}_{\text{enkf},t}^{(\theta)}(y_{0:t}|\tilde{x}_{0:t}^{1:N})$ for proposed parameter values that are likely to be rejected. Hence, we propose to replace the standard Metropolis-Hastings move in step 3 of Algorithm 6 with a delayed acceptance (DA) move (e.g. christen2005; Goli15; Banterle19).
@@
For a given particle $\theta^{i}$, Stage One of the DA scheme proposes $\theta^{*}\sim q(\theta^{i},\cdot)$, computes $\widehat{p}_{\textrm{knn},t}^{(\theta^{*})}(y_{0:t})$ and the screening acceptance probability
$\alpha_{1}\left(\theta^{i},\theta^{*}\right)=1\wedge\frac{\pi(\theta^{*})\widehat{p}_{\textrm{knn},t}^{(\theta^{*})}(y_{0:t})}{\pi(\theta^{i})\widehat{p}_{\textrm{knn},t}^{(\theta^{i})}(y_{0:t})}\times\frac{q(\theta^{*},\theta^{i})}{q(\theta^{i},\theta^{*})}.$ (19)
If this screening step is successful, Stage Two of the DA scheme constructs $\widehat{p}_{\text{enkf},t}^{(\theta^{*})}(y_{0:t}|\tilde{x}_{0:t}^{*,1:N})$ and the Stage Two acceptance probability
$\alpha_{2|1}(\theta^{i},\theta^{*})$ $=1\wedge\frac{\widehat{p}_{\text{enkf},t}^{(\theta^{*})}(y_{0:t}|\tilde{x}_{0:t}^{*,1:N})}{\widehat{p}_{\text{enkf},t}^{(\theta^{i})}(y_{0:t}|\tilde{x}_{0:t}^{i,1:N})}\times\frac{\widehat{p}_{\textrm{knn},t}^{(\theta^{i})}(y_{0:t})}{\widehat{p}_{\textrm{knn},t}^{(\theta^{*})}(y_{0:t})}.$ (20)
Thus the overall acceptance probability for the scheme is $\alpha_{1}\alpha_{2|1}$ and standard arguments show that iterating the DA scheme defines a Markov chain that is reversible with respect to the (extended) target $\widehat{\pi}_{\textrm{enkf},t}(\theta,\tilde{x}_{0:t}^{1:N},x_{0:t}^{1:N}|y_{0:t})$.

### v2

@@ 3.2 Ensemble MCMC
where $\psi_{T}^{(\theta)}(\tilde{x}_{0:T}^{1:N},x_{0:T}^{1:N})$ denotes the joint density of the entire ensemble $\{\tilde{x}_{0:T}^{1:N},x_{0:T}^{1:N}\}$. Although (14) is typically intractable, it can be sampled using MCMC, and the parameter samples retained to give (dependent) draws from $\widehat{\pi}_{\textrm{enkf},T}(\theta|y_{0:T})$, the $\theta$-marginal of (14). The resulting approach, termed ensemble MCMC (eMCMC) is detailed in Drovandi et al. (2022) (see also Stroud et al., 2018; Katzfuss et al., 2019). In brief, at iteration $k$, a new $\theta^{*}$ is proposed from a kernel $q(\theta^{k-1},\cdot)$. Then, sampling of the corresponding ensemble $\{\tilde{x}_{0:T}^{*,1:N},x_{0:T}^{*,1:N}\}$ from $\psi_{T}^{(\theta^{*})}(\cdot,\cdot)$ and evaluation of the observed data likelihood $\widehat{p}_{\textrm{enkf},T}^{(\theta^{*})}(y_{0:T}|\tilde{x}_{0:T}^{*,1:N})$, is achieved through recursive application of Algorithm 3. The parameter $\theta^{*}$ is retained with probability
@@
When applied sequentially, eMCMC is wasteful in terms of both storage and computational costs; upon receipt of a new observation (say $y_{T+1}$), previous parameter samples are discarded and eMCMC is run from scratch. Nevertheless, eMCMC is a key component of the resample-move step described in Section 4.
@@ 3.3 State augmentation
The EnKF can be directly modified to simultaneously estimate both hidden states and static parameters by augmenting the state vector $x_{t}$ to incorporate $\theta$ (see e.g. Evensen, 2009; Mitchell and Arnold, 2021). We write $z_{t}=(x_{t}^{\prime},\theta_{t}^{\prime})^{\prime}$ with $x_{t}$ evolving according to $p_{t}^{(\theta)}(\cdot|x_{t-1})$ and $\theta_{t}=\theta_{t-1}=\theta$. The observation model becomes $f_{t}^{(\theta)}(y_{t}|z_{t})=\mathcal{N}(y_{t};H_{z}z_{t},R)$ with $H_{z}=(H,0_{d_{y}\times d_{\theta}})$, where $0_{d_{y}\times d_{\theta}}$ is a $d_{y}\times d_{\theta}$ matrix of zeros and $H$ is not a function of $\theta$.
@@
To alleviate under-dispersion in the parameter ensemble, for example when small ensemble sizes are considered (e.g. Ruckstuhl and Janjić, 2018), the parameter vector can be allowed to dynamically evolve, for example using the artificial evolution kernel of Liu and West (2001), which gives
@@
where $\bar{\theta}_{t-1}$ and $V_{t-1}$ are the sample mean and variance of $\theta_{t-1}^{1},\ldots,\theta_{t-1}^{N}$. Setting $a=\sqrt{1-h^{2}}$ corrects for over-dispersion; the practitioner can then choose the smoothing parameter $h$, for example using the discounting method of Liu and West (2001). Step $t$ of the augmented EnKF is summarised by Algorithm 4.
@@ 3.4 Particle ensemble Kalman filter (PEnKF)
The augmented EnKF makes the strong assumption that the parameters and observations follow a linear Gaussian structure. An alternative approach for dealing with the static parameters is to marginalise out the hidden states in a sequential fashion. Recall the factorisation of $\pi_{t}(\theta|y_{0:t})$ in (5) and the assumption of a weighted sample of size $M$, $\{\theta^{1:M},u_{t-1}^{1:M}\}$, from $\pi_{t-1}(\theta|y_{0:t-1})$. Upon receipt of $y_{t}$, and in the context of time-varying parameters, Katzfuss et al. (2019) (see also Frei and Künsch, 2011, among others) propose to use the observed-data likelihood under the EnKF to reweight parameter particles. We adapt the resulting particle ensemble Kalman filter (PEnKF) to our problem as follows.
@@ 4 Nested ensemble Kalman filter (NEnKF)
From the perspective of static parameter filtering, the EnKF-based methods described in Section 3 attempt to mitigate particle degeneracy via an artificial evolution kernel over the static parameters. This can be sensitive to the choice of tuning parameters (Vieira and Wilkinson, 2016). The SMC2 algorithm described in Section 2 provides a more principled approach to alleviating particle degeneracy: if triggered at some time $t$, a resample-move step mutates parameter particles through a (pseudo-marginal) Metropolis-Hastings kernel that has $\pi_{t}(\theta|y_{0:t})$ as its invariant distribution. However, both the weighting and resample-move step in SMC2 require running a particle filter over states, which can require a computationally prohibitive number of state particles $N$ in high-dimensional settings. For a specific example with $d_{x}=d_{y}=:n$, Katzfuss et al. (2019) show that the variance of the logarithm of observed-data likelihood under the EnKF, $\textrm{Var}[\log\widehat{p}_{\textrm{enkf},T}^{(\theta)}(y_{0:T}|\tilde{x}_{0:T}^{1:N})]$, can be controlled by scaling $N$ linearly with $n$, whereas $N$ should be scaled exponentially with $n$ when controlling $\textrm{Var}[\log\widehat{p}_{T}^{(\theta)}(y_{0:T}|x_{0:T}^{1:N},a_{0:T-1}^{1:N})]$, computed using a particle filter. When comparing the efficiency of ensemble MCMC and particle MCMC across several examples, Drovandi et al. (2022) find that the former can tolerate a much smaller value of $N$ compared to the latter.
@@ 4.1 NEnKF algorithm
$\rho_{\textrm{enkf}}:=\frac{\pi(\theta^{*})}{\pi(\theta^{i})}\times\frac{\widehat{p}_{\textrm{enkf},t}^{(\theta^{*})}(y_{0:t}|\tilde{x}_{0:t}^{*,1:N})}{\widehat{p}_{\textrm{enkf},t}^{(\theta^{i})}(y_{0:t}|\tilde{x}_{0:t}^{i,1:N})}\times\frac{q(\theta^{*},\theta^{i})}{q(\theta^{i},\theta^{*})}.$ (18)
@@
   2. Resample. Normalise the weights to give $u_{t}^{i}=\tilde{u}_{t}^{i}/\sum_{k=1}^{M}\tilde{u}_{t}^{k}$. Sample an index $b_{i}$ from $\{1,\ldots,M\}$ with probabilities $u_{t}^{1:M}$. Put $\{\theta^{i},u_{t}^{i}\}:=\{\theta^{b_{i}},1/M\}$ and $x_{t}^{i,1:N}:=x_{t}^{b_{i},1:N}$.
   3. Move / mutate. Propose $\theta^{*}\sim q(\theta^{i},\cdot)$. Perform iterations $1,\ldots,t$ of Algorithm 3 to obtain $\widehat{p}_{\textrm{enkf},t}^{(\theta^{*})}(y_{0:t}|\tilde{x}_{0:t}^{*,1:N})$ and $x_{t}^{*,1:N}$. With probability ${\alpha}_{\textrm{enkf}}(\theta^{*}|\theta^{i})=1\wedge\rho_{\textrm{enkf}}$, where $\rho_{\textrm{enkf}}$ is given by (18), put $\theta^{i}=\theta^{*}$, $\widehat{p}_{\textrm{enkf},t}^{(\theta^{i})}(y_{0:t}|\tilde{x}_{0:t}^{i,1:N})=\widehat{p}_{\textrm{enkf},t}^{(\theta^{*})}(y_{0:t}|\tilde{x}_{0:t}^{*,1:N})$ and $x_{t}^{i,1:N}=x_{t}^{*,1:N}$.
@@
Following the approach used in SMC2, we also consider a mechanism for dynamically updating the number of ensemble members $N$. After completion of the mutation step, if updating $N$ is triggered then for each particle $\theta^{i}$, ensemble members $x_{0:t}^{i,1:N}$ are replaced by $\check{x}_{0:t}^{i,1:\check{N}}$ using the generalised importance sampling strategy of Del Moral et al. (2006), with an artificial backward kernel chosen to give an incremental weight of 1 (Botha et al., 2023b). Hence, for each particle $\theta^{i}$, a new ensemble $\check{x}_{0:t}^{i,1:\check{N}}$ is generated through repeated application of Algorithm 3. We adapt the method of Duan and Fulop (2015) to our context, by updating $N$ if $\hat{\sigma}_{N}^{2}>1.5$, where $\hat{\sigma}_{N}^{2}$ is an estimate of the variance of $\log\widehat{p}_{\text{enkf},t}^{(\bar{\theta})}(y_{0:t}|\tilde{x}_{0:t}^{1:N})$, based on, say, $r$ runs of the EnKF with $\theta$ fixed at an estimate of some central posterior value (e.g. the posterior mean). When the updating of $N$ is triggered, we set $\check{N}=\hat{\sigma}_{N}^{2}N$. We refer the reader to Botha et al. (2023b) for further discussion, noting their critique that, in the SMC2 setting, this approach can be sensitive to the initial value of $N$. However, and as discussed at the beginning of this section, $\widehat{p}_{\text{enkf},t}^{(\theta)}(y_{0:t}|\tilde{x}_{0:t}^{1:N})$ is likely to be less sensitive to $N$ than the corresponding particle filter estimator.
@@ 4.2 Resample-move: delayed-acceptance kernel
As with SMC2, the most computationally expensive step of the NEnKF algorithm is the resample-move step; if executed at time $t$, the cost is $\mathcal{O}(tMN)$. We, therefore, aim to avoid expensive calculation of the likelihood $\widehat{p}_{\text{enkf},t}^{(\theta)}(y_{0:t}|\tilde{x}_{0:t}^{1:N})$ for proposed parameter values that are likely to be rejected. Hence, we propose to replace the standard Metropolis-Hastings move in step 3 of Algorithm 6 with a delayed acceptance (DA) move (e.g. Christen and Fox, 2005; Golightly et al., 2015; Banterle et al., 2019).
@@
For a given particle $\theta^{i}$, Stage One of the DA scheme proposes $\theta^{*}\sim q(\theta^{i},\cdot)$, computes $\widehat{p}_{\textrm{knn},t}^{(\theta^{*})}(y_{0:t})$ and the screening acceptance probability $\alpha_{1}\left(\theta^{i},\theta^{*}\right)=1\wedge\rho_{1}$, where
$\rho_{1}:=\frac{\pi(\theta^{*})\widehat{p}_{\textrm{knn},t}^{(\theta^{*})}(y_{0:t})}{\pi(\theta^{i})\widehat{p}_{\textrm{knn},t}^{(\theta^{i})}(y_{0:t})}\times\frac{q(\theta^{*},\theta^{i})}{q(\theta^{i},\theta^{*})}.$ (19)
If this screening step is successful, Stage Two of the DA scheme constructs $\widehat{p}_{\text{enkf},t}^{(\theta^{*})}(y_{0:t}|\tilde{x}_{0:t}^{*,1:N})$ and the Stage Two acceptance probability $\alpha_{2|1}(\theta^{i},\theta^{*})=1\wedge\rho_{2|1}$, where
$\rho_{2|1}:=\frac{\widehat{p}_{\text{enkf},t}^{(\theta^{*})}(y_{0:t}|\tilde{x}_{0:t}^{*,1:N})}{\widehat{p}_{\text{enkf},t}^{(\theta^{i})}(y_{0:t}|\tilde{x}_{0:t}^{i,1:N})}\times\frac{\widehat{p}_{\textrm{knn},t}^{(\theta^{i})}(y_{0:t})}{\widehat{p}_{\textrm{knn},t}^{(\theta^{*})}(y_{0:t})}.$ (20)
The overall acceptance probability for the scheme is $\alpha_{1}\alpha_{2|1}$ and standard arguments show that iterating the DA scheme defines a Markov chain that is reversible with respect to the (extended) target $\widehat{\pi}_{\textrm{enkf},t}(\theta,\tilde{x}_{0:t}^{1:N},x_{0:t}^{1:N}|y_{0:t})$.
By design (to ensure detailed balance), $\rho_{1}\rho_{2|1}=\rho_{\textrm{enkf}}$. Hence $\alpha_{1}\alpha_{2|1}\leq\alpha_{\textrm{enkf}}$, since for any $a,b\geq 0$, $\{1\wedge a\}\{1\wedge b\}\leq 1\wedge ab$. Thus, although the delayed acceptance step can substantially reduce the number of computationally intensive and potentially wasteful estimations of the EnKF likelihood, it also reduces the expected amount of movement. Intuitively, this reduction will be small when $\rho_{2|1}\approx 1$. More formally, we may obtain a bound on the worst possible behaviour:
$\alpha_{1}=1\wedge\left\{\rho_{\textrm{enkf}}\times\frac{1}{\rho_{2|1}}\right\}\geq\left\{1\wedge\rho_{\textrm{enkf}}\right\}\times\left\{1\wedge\frac{1}{\rho_{2|1}}\right\}=\alpha_{\textrm{enkf}}\times\left\{1\wedge\frac{1}{\rho_{2|1}}\right\}.$
This provides a lower bound for $\alpha_{\textrm{da}}$, the overall acceptance probability of the delayed acceptance move:
$\alpha_{\textrm{da}}=\alpha_{1}\alpha_{2|1}\geq\alpha_{\textrm{enkf}}\times\left\{1\wedge\frac{1}{\rho_{2|1}}\right\}\times\left\{1\wedge\rho_{2|1}\right\}=\alpha_{\textrm{enkf}}\times\exp\{-|\log\rho_{2|1}|\}.$
In general, $|\log\rho_{2|1}|$ is small when
$\widehat{p}_{\textrm{knn},t}^{(\theta^{i})}(y_{0:t})\approx\widehat{p}_{\text{enkf},t}^{(\theta^{i})}(y_{0:t}|\tilde{x}_{0:t}^{i,1:N})~~~\mbox{and}~~~\widehat{p}_{\textrm{knn},t}^{(\theta^{*})}(y_{0:t})\approx\widehat{p}_{\text{enkf},t}^{(\theta^{*})}(y_{0:t}|\tilde{x}_{0:t}^{*,1:N}).$
With our inverse-distance weighting scheme, the first approximate equality is, in fact, an equality. A necessary condition for the second approximate equality is that both $\widehat{p}_{\text{enkf},t}^{(\theta^{*})}(y_{0:t}|\tilde{x}_{0:t}^{*,1:N})$ and $\widehat{p}_{\textrm{knn},t}^{(\theta^{*})}(y_{0:t})$ are close to $E^{\theta^{*}}_{t}:=\text{E}\left[\widehat{p}_{\text{enkf},t}^{(\theta^{*})}(y_{0:t}|\tilde{X}_{0:t}^{*,1:N})\right]$. Keeping $\hat{\sigma}_{N}^{2}<1.5$ helps ensure closeness for the former, most of the time; for example, if $\widehat{p}_{\text{enkf},t}^{(\theta^{*})}(y_{0:t}|\tilde{X}_{0:t}^{*,1:N})$ has a lognormal distribution then $\mathbb{P}\left(|\log\widehat{p}_{\text{enkf},t}^{(\theta^{*})}(y_{0:t}|\tilde{x}_{0:t}^{*,1:N})-\log E^{\theta^{*}}_{t}|<1\right)>1/2$. The $k$-nearest neighbour approximation averages $k$ noise terms, so the variability in this approximation to $\log E^{\theta^{*}}_{t}$ will typically be substantially less than $1.5$. Thus, we expect the discrepancy here to be driven by how well the weighted average of $E^{\theta^{[i]}}_{t},\dots,E^{\theta^{[k]}}_{t}$ approximates $E^{\theta^{*}}_{t}$. This will hold provided $\log E^{\theta}_{t}$ varies sufficiently slowly with $\theta$ compared with the spacing of the $N$ points. Thus, we would expect the $k$-nearest neighbour strategy to fail when the parameter space is high-dimensional or when the log posterior experiences high-amplitude oscillations over short distances.